OpenStudy (anonymous):

Let A = [cos θ -sin θ] [sin θ cos θ] Show that A^n = [cos(nθ) -sin(nθ)] [sin(nθ) cos(nθ)]. Is this plausible? Explain. (Note: Those are matrices mentioned above.)

OpenStudy (kinggeorge):

It certainly seems plausible. Try the case where n=2 first, and then you should be able to prove it for all n.

OpenStudy (anonymous):

I tried n = 2 as well and it works as well. How can I explain this if it is plausible?

OpenStudy (kinggeorge):

When you tried n=2, simply erase every 2, and rewrite n in it's place. This should be sufficient to show that it for A^n.

OpenStudy (kinggeorge):

A better way to explain in that I just noticed, would be to say that this is a rotation matrix. It rotates a vector by \(\theta\) in the x-y plane. Thus, if you squared it, you would apply the rotation twice, and would rotate by \(2\theta\). Then, you can say that if you applied the transformation n-times (A^n), you would be rotating by \(n\theta\). So you should have \(n\theta\) in place of \(\theta\).

OpenStudy (anonymous):

I get it now. Thanks for the help. :)

OpenStudy (kinggeorge):

Sorry that took so long.

OpenStudy (anonymous):

No prob it's alright. It was a bit challenging lol.